# Pretraining with Random Noise for Fast and Robust Learning without Weight Transport

### v1

@@ 4.2 Pretraining random noise enables fast learning during subsequent data training
Figure 2: Effect of random noise training on subsequent data training. (a) Design of the MNIST classification task to investigate the effect of random training. (b) Test accuracy during the training process, where the inset demonstrates the convergence speed of each training method, calculated by the area under the curve (AUC) of the test accuracy. (c) Alignment angle between weights and synaptic feedback across random training and data training. (d) Trajectory of weights ($\mathbf{W}_{1}$) toward synaptic feedback matrix ($\mathbf{B}_{1}$) in latent space obtained by PCA (principal component analysis) for random and data training. (e) Distance between the weights ($\mathbf{W}_{1}$) and the synaptic feedback matrix ($\mathbf{B}_{1}$). (f) Order dependence of the trajectory of the weights ($\mathbf{W}_{1}$). (g) Distance between the weights ($\mathbf{W}_{1}$) and the synaptic feedback matrix ($\mathbf{B}_{1}$) for different orders of random and data trainings.
@@
Next, we compared networks with and without random training in terms of subsequent data training outcomes (Figure 2a). We trained the networks with the MNIST dataset [52], a widely used dataset for image classification benchmarking, finding that a randomly trained network can learn the data more quickly compared to a network without random training (Figure 2b). To quantify the speed of learning, we calculated the area under the curve (AUC) of the test accuracy and found that the convergence of the randomly trained network is significantly faster than that in the network without random training (Figure 2b, inset, untrained network vs. randomly trained network, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). Notably, the convergence speed of the randomly trained network appeared comparable to that of the network trained with backpropagation (Figure 2b, inset, randomly trained network vs. backpropagation, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). We also observed that the weight alignment gap between untrained and randomly trained networks is maintained during data training (Figure 2c). As a result, at the end of the data training step, the alignment angle of the randomly trained network was significantly smaller than that of the untrained network (Figure 2c, alignment angle in an untrained network vs. a randomly trained network, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that a combination of random pretraining and subsequent data training can enhance the weight alignment, which leads to more precise error teaching.
@@
To understand the weight update dynamics by random and data training, we visualized the trajectory of weights in latent space as obtained by a principal component analysis (PCA) [53] (Figure 2d). We conducted PCA on the weights of the last layer ($\mathbf{W}_{1}$) for the random and data training conditions. First, we confirmed that in both random and data training, the weights become closer to synaptic feedback (Figure 2e, untrained vs. randomly trained network, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$; untrained vs. data trained network, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$; randomly trained network vs. data trained network, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). Notably, we observed that the updated trajectory of weights by random training and data training have different directions in the principal component space and that the effects of random training depend on the order of the random and data training (Figure 2f, g) — the enhancement of weight alignment was more significant when data training was performed after random training compared to when training is done in a reversed order. Particularly, when we trained the network with data first, subsequent random training could not move the weights (Figure 2g, data trained vs. data and random trained, $n_{\text{Net}}=10$, t-test, NS, $P=0.999$); thus, the weights did not become closer to synaptic feedback (Figure 2g, random trained vs. random and data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that weight alignment by random noise pretraining cannot be replaced by data training and that it is crucial to perform random training prior to data training.
@@ 4.3 Pre-regularization by random noise training enables robust generalization
Figure 3: Pre-regularization by random noise training enhances generalization (a) Untrained network and pre-regularized network through random noise training. (b) Distribution of the readout probability. (c) Singular value spectrum of the forward weights. (d) Effective rank of forward weights during random noise training. (e) Generalization error gap between the training error and test error (training set size: 1600, network depth: 3). (f) Generalization error for various training set sizes (network depth: 3). (g) Effective dimensionality of the Gram matrix, the cosine similarity of feature vectors across neurons (training set size: 1600, network depth: 3). (h) Effective dimensionality of the Gram matrix for various network depths (training set size: 1600).
Next, we compared the difference between an untrained network and a randomly trained network in terms of their activation and weight (Figure 3a). First, we found that the readout probability of the untrained network is distributed over a wide range (Figure 3b, left, readout probability vs. chance level, $n=10,000$, one-sample t-test, ${}^{*}P<0.001$), whereas that of the randomly trained network is well regularized, close to the chance level (Figure 3b, right, readout probability vs. chance level, $n=10,000$, one-sample t-test, NS, $P=0.371$). We also observed that the singular value spectrum of forward weights changes significantly by random training (Figure 3c) such that a small portion of singular values become dominant in the randomly trained network. To measure the effective dimensionality of the weights quantitatively, we used the effective rank of the weights.
Definition. Given a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is decomposed into $\mathbf{A}=U\Sigma V^{T}$ by singular value decomposition (SVD), the singular values are $\{\sigma_{i}\}_{i}^{min(m,n)}$ sorted in a descending order. The effective rank $\rho$ is defined as the Shannon entropy of the normalized singular values, $\rho=-\sum_{i}{\bar{\sigma}_{i}}\log{\bar{\sigma_{i}}}$, where $\bar{\sigma}_{i}=\sigma_{i}/\sum_{i}\sigma_{i}$. Without loss of generality, we used the effective rank as the exponential of $\rho$ [54].
@@
To test the generalization ability of the network, we measured the gap between the training error and the test error during subsequent data training. We observed that the generalization error was noticeably lower in a randomly trained network compared to an untrained network (Figure 3e) and that this tendency was maintained with variations of the training set size (Figure 3f, random and data trained vs. data trained,$n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that pre-regularization by random training can enable robust generalization during subsequent data training.
Next, we compared the representation of learned features in an untrained network and a randomly trained network. We used the Gram matrix, defined as the cosine similarity of feature vectors across neurons. Notably, we found that the effective rank of the Gram matrix was significantly lower in a randomly trained network compared to an untrained network after subsequent data training (Figure 3g, random and data trained vs. data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$) and that this tendency was maintained regardless of the network depth (Figure 3h, random and data trained vs. data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). This finding suggests that pre-regularization by random training can enable networks to learn simpler solutions, leading to better generalization performance during subsequent data training.
Figure 4: Robust generalization of “out-of-distribution“ tasks in randomly trained networks. (a) Training in-distribution data (MNIST) in untrained and randomly trained networks. (b) Out-of-distribution generalization tests on transformed MNIST. (c) Out-of-distribution generalization tests on USPS dataset.
We also tested the generalization performance of the networks for “out-of-distribution” tasks by training the network with the MNIST dataset and testing it with various out-of-distribution tasks (Figure 4a). First, we generated a MNIST dataset of translated, rotated, and scaled images and then used these images as out-of-distribution tasks (Figure 4b, left). We observed that a randomly trained network showed significantly higher test accuracy on out-of-distribution tasks than an untrained network (Figure 4b, right, random and data trained vs. data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). We also observed that the randomly trained network showed higher test accuracy on the USPS dataset, which is a widely used benchmark dataset for out-of-distribution tasks (Figure 4c, random and data trained vs. data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that pre-regularization by random training enables robust out-of-distribution generalization during subsequent data training.
@@ 4.4 Task-agnostic fast learning for various tasks by a network pretrained with random noise
Figure 5: Task-agnostic fast learning for various tasks in randomly trained networks. (a) Three tasks used to test the task-agnostic property of random training, showing the meta-loss during the random training process. The meta-loss is calculated from the sum of the losses measured during adaption to each task. (b) Trajectory of weights in the latent space for adaptation to each task of an untrained network and a randomly trained network. (c) Adaption to each task of an untrained network and a randomly trained network.
Lastly, we examined whether random training is task-agnostic. We compared the task adaption capacity of an untrained network and a randomly trained network on three tasks: (1) MNIST classification [52], (2) Fashion-MNIST [55], (3) Kuzushiji-MNIST [56] (Figure 5a, left). To measure the ability of fast adaption to various tasks quantitatively, we computed the meta-loss, as suggested in a previous study of meta-learning.
Definition. Given the task distribution $\mathcal{T}\in\{\mathcal{T}_{i}\}_{i}^{n}$, the meta-loss of network $f_{\theta}$ is defined as $\mathcal{L}_{\text{meta}}(\theta)=\sum_{\mathcal{T}_{i}\in\mathcal{T}}\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i}^{{}^{\prime}})$, where $\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i}^{{}^{\prime}})$ denotes the loss of the task $\mathcal{T}_{i}$ and $\theta_{i}^{{}^{\prime}}$ is the adapted parameter for $\mathcal{T}_{i}$ [57].
We observed that the meta-loss decreased gradually during the random training process (Figure 5a, right). Considering that the training was solely performed with random inputs and labels on the three tasks to measure the meta-loss, this result suggests that networks can learn how to adapt to various tasks without any task-specific data. Next, we trained the untrained networks and random noise trained networks on each task separately. We conducted principal component analysis (PCA) on the weights of the last layer ($\mathbf{W}_{2}$) to visualize the trajectory of weights in latent space during the adaption to each task. We observed that the trajectory of weights during random noise training moves closer to synaptic feedback ($\mathbf{B}_{2}$), which makes the adaption to each task more efficient (Figure 5b). This suggests that random training is task-agnostic but provides efficient and fast learning in subsequent learning (Figure 5b). Lastly, we compared the adaption to each task in an untrained network and a randomly trained network. We observed that the randomly trained network showed remarkably fast adaption to each task compared to the untrained network (Figure 5c, random and data trained vs. data trained, $n_{\text{Net}}=10$, t-test, ${}^{*}P<0.001$). These results highlight the task-agnostic property of random training, which enables networks quickly to adapt to various tasks.
@@ 5 Discussion
We showed that random noise training enables neural networks to learn quickly and robustly without weight transport. This finding bridges the gap between a biologically plausible learning mechanism and the conventional backpropagation algorithm, as symmetry of forward and backward weights can easily be achieved by random noise training. Moreover, the results here provide new insight into the advantage of random training as a means of pre-regularization of a network for robust generalization.
Insights into developmental neuroscience. Unlike artificial neural networks, the brain is ready to learn before encountering data. In the early developmental stage before eye-opening, spontaneous random activity emerges in the brain, which is considered essential for a normal development of early circuits. However, the functional advantage of learning from random noise before external sensory inputs remains unclear. Our study provides a plausible scenario that the brain utilizes spontaneous random activity to pre-align the synaptic weights for error learning and pre-regularization of synaptic connections for robust generalization. Specifically, we showed that random training reduces the effective dimensionality of the weights, which can be considered as a form of pruning, as previous neuroscience studies reported that the brain’s synaptic connections are pruned substantially during development, particularly dependent on spontaneous activity [27, 28, 29, 30, 31]. Despite the fact that the present study is based on model neural networks, the results here are consistent with a range of experimental findings in developmental neuroscience.
@@
Error-backpropagation without weight transport. Understanding how the brain can efficiently learn is a collective goal among neuroscience and artificial intelligence researchers. Early works by neuroscientists have suggested several naive learning rules, such as Hebbian learning [58] and spike-timing-dependent plasticity (STDP) [59, 60, 61]. Although these rules have been experimentally observed and are thus biologically plausible, they are not sufficient to explain the brain’s remarkable learning ability thoroughly [62, 63]. On the other hand, while the backpropagation algorithms used in artificial neural networks have shown impressive learning capabilities, they are considered biologically implausible due to the weight transport problem [36]. Our results provide a new perspective on this issue, bridging the gap between these training rules. We showed that symmetry among forward and backward weights, which is necessary to back-propagate errors, can be readily developed by learning random noise, similar to that during the brain’s prenatal stage. Our findings suggest a probable scenario for significantly narrowing the performance gap between previously suggested biologically plausible learning rules and backpropagation.
@@
Pre-regularization for robust generalization. We suggest a task- and model-agnostic pretraining strategy that involves simply training the network with random noise. Notably, our results here show that random noise training can enhance the learning efficiency and generalization ability of the network, for which various tricks and techniques have been proposed to improve. We found that pretraining on random noise reduces the effective dimensionality of the weights, which leads to the learning of low-rank solutions for tasks. Previous studies on generalization have shown that the low-rank bias of neural networks is crucial for their generalization ability, as we confirmed here [64, 65, 66, 67, 68]. Also, our results showed that the complexity of the representation is significantly reduced in a network pretrained with random noise, which might be a type of information bottleneck [69, 70, 71, 72, 73]. Also, the refinement of readout activity to the chance level suggests that random noise training can transform the loss function of the network to a smoother and more sensitive form for learning [74, 75, 76]. It is also important to note that the simple strategy of random noise training can have various effects on the network’s learning dynamics that previous machine learning techniques have aimed to achieve. This may be a possible strategy employed by the brain to achieve notable generalization ability. At the same time, it suggests a new form of pretraining strategy for artificial neural networks.
6 Limitations and Broader Impacts
Limitations. Although our study provides a new perspective on the role of spontaneous random activity in the brain, there are several limitations to consider. We used a fully connected feedforward neural network. It shares some similarities with biological neural networks, but they are distinct in terms of structure and function. Also, we used a feedback alignment algorithm to train the networks. As it does not utilize weight transport, further experimental studies are needed to confirm our results in actual biological systems.

### v2

@@ 4.2 Pretraining random noise enables fast learning during subsequent data training
Figure 2: Effect of random noise pretraining on subsequent data training. (a) Design of the MNIST classification task to investigate the effect of random training. (b) Test accuracy during the training process, where the inset demonstrates the convergence speed of each training method, calculated by the AUC of the test accuracy. (c) Alignment angle between weights and synaptic feedback across random training and data training. (d) Trajectory of weights ($\mathbf{W}_{1}$) toward synaptic feedback ($\mathbf{B}_{1}$) in latent space obtained by PCA for random and data training. (e) Distance between the weights ($\mathbf{W}_{1}$) and the synaptic feedback ($\mathbf{B}_{1}$). (f) Order dependence of the trajectory of the weights ($\mathbf{W}_{1}$). (g) Distance between the weights ($\mathbf{W}_{1}$) and the synaptic feedback ($\mathbf{B}_{1}$) for different orders of random and data trainings.
@@
Next, we compared networks with and without random pretraining in terms of subsequent data training outcomes (Figure 2a). We trained the networks using a subset of the MNIST dataset [52], a widely used benchmark for image classification. We found that a randomly pretrained network learns the data more quickly and achieves higher accuracy compared to a network that is not randomly pretrained (Figure 2b). To quantify the speed of learning, we calculated the area under the curve (AUC) of the test accuracy and found that the convergence of the randomly pretrained network is significantly faster than that in the network without random pretraining (Figure 2b, inset, w/o vs. w/ random pretraining (FA), $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). Notably, the convergence speed of the randomly pretrained network appeared comparable to that of the network trained with backpropagation (Figure 2b, inset, w/ random pretraining (FA) vs. BP, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). We also observed that the weight alignment gap between untrained and randomly pretrained networks is maintained during data training (Figure 2c). As a result, at the end of the data training step, the alignment angle of the randomly trained network was significantly smaller than that of the untrained network (Figure 2c, w/o vs. w/ random pretraining, $n=100$, t-test, ${}^{*}P<0.001$). This result suggests that a combination of random pretraining and subsequent data training can enhance the weight alignment, which leads to more precise error teaching.
@@
To understand the weight update dynamics by random and data training, we visualized the trajectory of weights in latent space as obtained by a principal component analysis (PCA) [53] (Figure 2d). We conducted PCA on the weights of the last layer ($\mathbf{W}_{1}$) for the random and data training conditions. First, we confirmed that in both random and data training, the weights become closer to synaptic feedback (Figure 2e, untrained vs. data training, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$; untrained vs. random training, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$; random vs. data training, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). Notably, we observed that the updated trajectory of weights by random training and data training have different directions in the principal component space and that the effects of random training depend on the order of the random and data training (Figure 2f, g) — the enhancement of weight alignment was more significant when data training was performed after random training (Figure 2g, random training vs. random and data training, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$) compared to when training is done in a reversed order (Figure 2g, data training vs. data and random training, $n_{\text{net}}=10$, t-test, NS, $P=0.999$). Particularly, when we trained the network with data first, subsequent random training could not move the weights; thus, the weights did not become closer to synaptic feedback. This result suggests that weight alignment by random noise pretraining cannot be replaced by data training and that it is crucial to perform random training prior to data training.
Figure 3: Comparison of model performance across different image datasets and network depths. (a-e) Final accuracy after convergence. Experiments were conducted with networks of varying depths on different tasks: (a) MNIST, (b) Fashion-MNIST, (c) CIFAR-10, (d) CIFAR-100, and (e) STL-10.
Table 1: Performance of the two-layer model for each dataset (MNIST, Fashion-MNIST, CIFAR-10, CIFAR-100, STL-10). Each performance value (%) is presented as the mean $\pm$ standard deviation from three trials. Extended results with various model depths can be found in the Appendix.
MNIST F-MNIST CIFAR-10 CIFAR-100 STL-10
BP $97.82\pm 0.03$ $88.87\pm 0.03$ $54.01\pm 0.20$ $24.55\pm 0.10$ $42.72\pm 0.20$
FA w/o $97.26\pm 0.07$ $87.47\pm 0.25$ $50.54\pm 0.22$ $20.17\pm 0.30$ $36.21\pm 0.91$
w/ $97.76\pm 0.07$ $88.26\pm 0.07$ $53.58\pm 0.12$ $24.45\pm 0.10$ $41.01\pm 0.16$
$\Delta$ACC $\blacktriangle 0.49\pm 0.06$ $\blacktriangle 0.79\pm 0.31$ $\blacktriangle 3.04\pm 0.20$ $\blacktriangle 4.28\pm 0.38$ $\blacktriangle 4.81\pm 0.86$
Next, we further investigated the model’s classification performance across various image datasets and network depths (Figure 3, Table 1). In earlier experiments with two-layer networks and MNIST, we showed that random noise pretraining enhances both the learning speed and accuracy of networks to levels comparable with backpropagation. We extended these experiments to networks of varying depths (Figure 3a) and confirmed that the benefits of random noise pretraining generalize to deeper networks. Additionally, we evaluated performance across various image datasets, including Fashion-MNIST [54] (Figure 3b), CIFAR-10 [55] (Figure 3c), CIFAR-100 [55] (Figure 3d), and STL-10 [56] (Figure 3e). We found that random noise pretraining significantly narrows the performance gap between feedback alignment and backpropagation across different datasets and depths. These results suggest that pretraining with random noise can serve as a general strategy for improving the performance of neural networks trained with feedback alignment algorithms, making it comparable to backpropagation.
@@ 4.3 Pre-regularization by random noise training enables robust generalization
Figure 4: Pre-regularization by random noise training enhances generalization (a) Untrained network and pre-regularized network through random noise training. (b) Distribution of the readout probability. (c) Singular value spectrum of the forward weights. (d) Effective rank of forward weights during random noise training. (e) Generalization error between the training error and test error (training set size: 1600, network depth: 3). (f) Generalization error for various training set sizes (network depth: 3). (g) Effective dimensionality of the Gram matrix, the cosine similarity of feature vectors across neurons (training set size: 1600, network depth: 3). (h) Effective dimensionality of the Gram matrix for various network depths (training set size: 1600).
Next, we compared the difference between an untrained network and a randomly trained network in terms of their activation and weight (Figure 4a). First, we found that the readout probability of the untrained network is distributed over a wide range (Figure 4b, top, untrained vs. chance level, $n=1000$, Wilcoxon signed-rank test, ${}^{*}P<0.001$), whereas that of the randomly trained network is well regularized, close to the chance level (Figure 4b, bottom, random noise trained vs. chance level, $n=1000$, Wilcoxon signed-rank test, NS, $P=0.096$). We also observed that the singular value spectrum of forward weights changes significantly by random training (Figure 4c) such that a small portion of singular values become dominant in the randomly trained network. To measure the effective dimensionality of the weights quantitatively, we used the effective rank of the weights.
Definition. Given a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ is decomposed into $\mathbf{A}=U\Sigma V^{T}$ by singular value decomposition (SVD), the singular values are $\{\sigma_{i}\}_{i=1}^{min(m,n)}$ sorted in descending order. The effective rank $\rho$ is defined as the Shannon entropy of the normalized singular values, $\rho=-\sum_{i}{\bar{\sigma}_{i}}\log{\bar{\sigma_{i}}}$, where $\bar{\sigma}_{i}=\sigma_{i}/\sum_{i}\sigma_{i}$. Without loss of generality, we used the effective rank as the exponential of $\rho$ [57].
@@
To test the generalization ability of the network, we measured the gap between the training error and the test error during subsequent data training. We observed that the generalization error was noticeably lower in a randomly pretrained network compared to a network trained solely on the data (Figure 4e) and that this tendency was maintained with variations of the training set size (Figure 4f, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that pre-regularization by random pretraining can enable robust generalization during subsequent data training.
Next, we compared the representation of learned features in a randomly pretrained network and a network trained without random pretraining. We used the Gram matrix, defined as the cosine similarity of feature vectors across neurons. Notably, we found that the effective rank of the Gram matrix was significantly lower in a randomly pretrained network compared to an untrained network after subsequent data training (Figure 4g, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$) and that this tendency was maintained regardless of the network depth (Figure 4h, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). This finding suggests that pre-regularization by random training can enable networks to learn simpler solutions, leading to better generalization performance during subsequent data training.
Figure 5: Robust generalization of "out-of-distribution" tasks in randomly pretrained networks. (a) Training in-distribution data (MNIST) in untrained and randomly pretrained networks. (b) Out-of-distribution generalization tests on transformed MNIST. (c) Out-of-distribution generalization tests on USPS dataset.
We also tested the generalization performance of the networks for "out-of-distribution" tasks by training the network with the MNIST dataset and testing it with various out-of-distribution tasks (Figure 5a). First, we generated a MNIST dataset of translated, rotated, and scaled images and then used these images as out-of-distribution tasks (Figure 5b, left). We observed that a randomly pretrained network showed significantly higher test accuracy on out-of-distribution tasks than a network trained without random pretraining (Figure 5b, right, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). We also observed that the randomly pretrained network showed higher test accuracy on the USPS dataset [58], which is a widely used benchmark dataset for out-of-distribution tasks (Figure 5c, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). This result suggests that pre-regularization by random pretraining enables robust out-of-distribution generalization during subsequent data training.
@@ 4.4 Task-agnostic fast learning for various tasks by a network pretrained with random noise
Figure 6: Task-agnostic fast learning for various tasks in randomly pretrained networks. (a) Three tasks used to test the task-agnostic property of random training, showing the meta-loss during the random training process. The meta-loss is calculated from the sum of the losses measured during adaptation to each task. (b) Trajectory of weights in the latent space for adaptation to each task of an untrained network and a randomly pretrained network. (c) Adaptation to each task of an untrained network and a randomly pretrained network.
Lastly, we examined whether random training is generally beneficial for various tasks. We compared the task adaptation capacity of an untrained network and a randomly pretrained network on three tasks: (1) MNIST classification [52], (2) Fashion-MNIST [54], (3) Kuzushiji-MNIST [59] (Figure 6a, left). To measure the ability of fast adaptation to various tasks quantitatively, we computed the meta-loss, as suggested in a previous study of meta-learning.
Definition. Given the task distribution $\mathcal{T}\in\{\mathcal{T}_{i}\}_{i=1}^{n}$, the meta-loss of network $f_{\theta}$ is defined as $\mathcal{L}_{\text{meta}}(\theta)=\sum_{\mathcal{T}_{i}\in\mathcal{T}}\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i}^{{}^{\prime}})$, where $\mathcal{L}_{\mathcal{T}_{i}}(\theta_{i}^{{}^{\prime}})$ denotes the loss of the task $\mathcal{T}_{i}$ and $\theta_{i}^{{}^{\prime}}$ is the adapted parameter for $\mathcal{T}_{i}$ [60].
We observed that the meta-loss decreased gradually during the random training process (Figure 6a, right). Considering that the training was solely performed with random inputs and labels on the three tasks to measure the meta-loss, this result suggests that networks can learn how to adapt to various tasks without any task-specific data. Next, we trained the untrained networks and randomly pretrained networks on each task separately. We conducted PCA on the weights of the last layer ($\mathbf{W}_{2}$) to visualize the trajectory of weights in latent space during the adaptation to each task. We observed that the trajectory of weights during random noise training moves closer to synaptic feedback ($\mathbf{B}_{2}$), which makes the adaptation to each task more efficient (Figure 6b). This suggests that random training is task-agnostic but provides efficient and fast learning in subsequent learning. Lastly, we compared the adaptation to each task in an untrained network and a randomly pretrained network. We observed that the randomly trained network showed remarkably fast adaptation to each task compared to the untrained network (Figure 6c, w/o vs. w/ random pretraining, $n_{\text{net}}=10$, t-test, ${}^{*}P<0.001$). These results highlight the task-agnostic property of random training, which enables networks to quickly adapt to various tasks.
@@ 5 Discussion
We showed that random noise pretraining enables neural networks to learn quickly and robustly without weight transport. This finding bridges the gap between a biologically plausible learning mechanism and the conventional backpropagation algorithm, as the symmetry of forward and backward weights can easily be achieved by random noise pretraining. Moreover, the results here provide new insight into the advantage of random training as a means of preconditioning a network for robust generalization.
@@
Error-backpropagation without weight transport. Early work in neuroscience identified basic learning rules, such as Hebbian learning [61] and spike-timing-dependent plasticity (STDP) [62, 63, 64]. Although these rules have been experimentally observed and are thus biologically plausible, they are not sufficient to explain the brain’s remarkable learning ability thoroughly [65, 66]. On the other hand, while the backpropagation algorithms used in artificial neural networks have shown impressive learning capabilities, they are considered biologically implausible due to the weight transport problem [23]. Our results provide a new perspective on this issue, bridging the gap between these training rules. We showed that symmetry among forward and backward weights, which is necessary to back-propagate errors, can be readily developed by learning random noise, similar to that during the brain’s prenatal stage. Our findings suggest a probable scenario for significantly narrowing the performance gap between previously suggested biologically plausible learning rules and backpropagation.
Recent studies explored noise as a biologically plausible mechanism to enhance learning efficiency without the need for weight transport. For instance, the weight mirror algorithm [36] uses noisy firing to align feedback weights with forward weights. Similarly, phaseless alignment learning [40] leverages layer-wise noise as an additional information carrier to achieve weight alignment. While these approaches have been reported to outperform traditional feedback alignment algorithm, merely incorporating noise into existing feedback alignment algorithms yields no improvement in learning performance [40]. In contrast, our results indicate that exposure to random noise "before encountering real data" significantly enhances vanilla feedback alignment. Unlike previous approaches, our strategy utilizes random noise for pre-conditioning the network, preparing it to learn more effectively. This aligns with biological observations that neural noise predominates in the early stage of brain development [67], even prior to exposure to external stimuli. It is important to note that our proposed method is not limited to the feedback alignment algorithm; pretraining with random noise could be beneficial for other algorithms, which we intend to explore in our follow-up studies.
@@
Pre-regularization for robust generalization. We suggest a task- and model-agnostic pretraining strategy that involves simply training the network with random noise. Notably, our results here show that random noise training can enhance the learning efficiency and generalization ability of the network, for which various tricks and techniques have been proposed to improve. We found that pretraining on random noise reduces the effective dimensionality of the weights, facilitating the learning of low-rank solutions for various tasks. Previous studies on generalization have shown that the low-rank bias of neural networks plays a crucial role in their generalization ability, a finding we have confirmed in this study [68, 69, 70, 71, 72]. Additionally, our results highlight that random noise pretraining functions as a form of meta-learning, enhancing the network’s ability to adapt rapidly to different tasks. In contrast to previous approaches that utilized data from diverse task distributions [60], our method achieves similar effects by merely training with noise. It is important to note that the straightforward strategy of random noise training can significantly influence the network’s learning dynamics - effects that previous machine learning techniques have sought to achieve. This approach may reflect a potential strategy employed by the brain to attain notable generalization capabilities. Furthermore, it suggests a novel pretraining strategy for artificial neural networks.
Insights into developmental neuroscience. Unlike artificial neural networks, the brain is ready to learn before encountering data. In the early developmental stage before eye-opening, spontaneous random activity emerges in the brain, which is considered essential for the normal development of early circuits [1, 2, 3, 4]. However, the functional advantage of learning from random noise before external sensory inputs remains unclear. Our study provides a plausible scenario that the brain utilizes spontaneous random activity to pre-align the synaptic weights for error learning and pre-regularization of synaptic connections for robust generalization. Specifically, we showed that random training reduces the effective dimensionality of the weights, which can be considered as a form of pruning, as previous neuroscience studies reported that the brain’s synaptic connections are pruned substantially during development, particularly dependent on spontaneous activity [4, 5, 6, 7, 8]. Despite the fact that the present study is based on model neural networks, the results here are consistent with a range of experimental findings in developmental neuroscience.
6 Broader impacts and limitations